# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 9 Preparatory Results
Let $A$ be a set, $a\in A$, and let $\rho,\tau,\omega$ be equivalence relations on $A$ such that $a/_{\rho}\subset a/_{\omega}$ and $a/_{\tau}\subset a/_{\omega}$. Then $a/_{\rho}\cup a/_{\tau}\subset a/_{\omega}$.
@@
There exist $b,c\in A$, such that $b\in a/_{\omega}\setminus a/_{\rho}$ and $c\in a/_{\omega}\setminus a/_{\tau}$. We now distinguish three cases.
Case 1: $(b,c)\in\rho$. Then by the definition of $b$, $(a,c)\not\in\rho$. By the definition of $c$, $(a,c)\not\in\tau$. Thence $(a,c)\not\in\rho\cup\tau$.
Case 2: $(b,c)\in\tau$. Then by the definition of $c$, $(a,b)\not\in\tau$. By the definition of $b$, $(a,b)\not\in\rho$. Consequently $(a,b)\not\in\rho\cup\tau$.
Case 3: $(b,c)\not\in\rho\cup\tau$. Then by the definition of $b$ and $c$, $(b,c)\in\omega$, and hence $(b,c)\in\omega\setminus(\rho\cup\tau)$.
In all three cases we have $a/_{\rho}\cup a/_{\tau}\subset a/_{\omega}$. ∎
@@
Let $A$ be a set, $a\in A$, and let $\rho,\tau,\omega$ be equivalence relations on $A$ such that $a/_{\rho}\cup a/_{\tau}=a/_{\omega}$. Then $a/_{\rho}=a/_{\omega}$ or $a/_{\tau}=a/_{\omega}$.
@@
By contradiction, assume that $a/_{\rho}\neq a/_{\omega}$ and $a/_{\tau}\neq a/_{\omega}$. Then $a/_{\rho}\subset a/_{\omega}$ and $a/_{\tau}\subset a/_{\omega}$. Consequently, by Proposition 9.79, $a/_{\rho}\cup a/_{\tau}\subset a/_{\omega}$. ∎
@@
[…]
@@
$(3\Rightarrow 4)$ $\mathop{\leftrightarrow}^{*}_{H}$ is a congrunce on the ground term algebra ${\cal T}(\Sigma)$, hence $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$.
@@
[…]
@@
[…]
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=size(E)+size(F)$. Let $W=ST\langle E\cup F\rangle$. We can decide in $O(n^{3})$ time whether for each $p\in W$,
@@
By Propositon 4.34, $s\mathop{\leftrightarrow}^{*}_{E}t$. Then
@@
We now introduce the notion of GTRS $R\langle E,F\rangle$ simulating the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant $a\in C\langle E\cup F,\emptyset\rangle$, then the concept of GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Then we describe the connection between the latter concept and the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and let $a\in C\langle E\cup F,\emptyset\rangle$. We say that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$ if Conditions 1 and 2 hold:
@@
    For every $b\in C\langle E\cup F,\emptyset\rangle$, if $R\langle{E\cup F},\emptyset\rangle$ reach $a$ from a proper extension of the constant $b$, then
    for every rewrite rule of the form $\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b\in R\langle{E\cup F},\emptyset\rangle$ with $m\in{\mathbb{N}}$,
[…]
    there exists $c\in C\langle E,F\rangle$ such that the rewrite rule $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c$ is in $R\langle E,F\rangle$.
Statement 10.89.
$c\subseteq b$ for $c,b$ appearing in Condition 2 in Definition 10.88.
Intuitively, in Condition 2 in Definition 10.88, the rewrite rule $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c$ in $R\langle E,F\rangle$ fits to the rewrite rule $\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b\in R\langle{E\cup F},\emptyset\rangle$.
Statement 10.90.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ is total, and let $a\in C\langle E\cup F,\emptyset\rangle$. GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$ if and only if $a\in C\langle E,F\rangle$.
Proof.
By direct inspection of Definition 4.29 and Definition 10.88. ∎
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. We say that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if for each constant $a\in C\langle E\cup F,\emptyset\rangle$, GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on $a$.
Statement 10.92.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ is total. GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant $a\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
1. $a\in C\langle E,F\rangle$.
2. $a\in C\langle F,E\rangle$ and GTRS $R\langle F,E\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on $a$.
Proof.
We have the satement by direct inspection of Definition 10.91 and Statement 10.90. ∎
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant $a\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
1. $a\in C\langle E,F\rangle$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
2. $a\in C\langle F,E\rangle$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
Proof.
By Definitions 10.88 and 10.91, and Lemma 6.70, and Definition 6.61, we have the lemma. ∎
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that $R\langle E,F\rangle$ is total. GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant $a\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
1. $a\in C\langle E,F\rangle$.
2. $a\in C\langle F,E\rangle$ and for each constant $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
Proof.
By Statement 10.92, Definition 10.88, Lemma 6.70, and Definition 6.61 we have the lemma. ∎
@@
The length of the first path is $1$, and it starts from the vertex $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
$f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$   and   #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$.
The length of the second path is $1$, and it starts from the vertex $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
$f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$   and   \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
@@
We now continue Example 3.18, Example 5.51, and Example 7.72. There does not exists a path of positive length in $AUX[E;F]$. There exists one cycle in $AUX[E;F]$. It is of length $1$, and it starts from the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
Observe that neither Condition 1 in nor Condition 2 in Lemma 10.93 holds for the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Thus by Lemma 10.93, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
Intuitively, assume that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a vertex $c\in C\langle E\cup F,\emptyset\rangle$ of $AUX[E;F]$. Consider a term $\delta[a_{1},\ldots,a_{k}]$ for some $\delta\in CON_{\Sigma,k}$, $a_{1},\ldots,a_{k}\in C\langle{E\cup F},\emptyset\rangle$ and a sequence of reductions
(i)   $\delta[a_{1},\ldots,a_{k}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}c$.
[…]
(ii)   $\delta[b_{1},\ldots,b_{k}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}d$
for some $d\in C\langle E,F\rangle$. We obtain (ii) following step by step the reductions in (i). For each reduction of (i), we apply a fitting rule at the same position in (ii). Then for each reduction step, for the two terms $s$ and $t$ obtained in (i) and (ii), respectively, each constant $b\in C\langle E,F\rangle$ appearing in $t$ corresponds to the constant $a\in C\langle E\cup F,\emptyset\rangle$ appearing at the same position in $s$. Moreover $b$, as congruence class, is subset of the corresponding constant $a$. Finally, constant $d$ corresponds to constant $c$, and $d\subseteq c$. Formally, we state and prove the following result.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant $c\in C\langle E\cup F,\emptyset\rangle$. Let Conditions (a) and (b) hold.
@@
    $\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-2}[a_{(n-2)1},\ldots,a_{(n-2)k_{n-2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[a_{11},\ldots,a_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}c$,
[…]
@@
[…]
$\delta_{n}[b_{n_{1}},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-2}[b_{(n-2)1},\ldots,b_{(n-2)k_{n-2}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\newline \delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}d$.

### v2

@@ 9 Preparatory Results
Let $A$ be a set, ${b}\in A$, and let $\rho,\tau,\omega$ be equivalence relations on $A$ such that ${b}/_{\rho}\subset{b}/_{\omega}$ and ${b}/_{\tau}\subset{b}/_{\omega}$. Then ${b}/_{\rho}\cup{b}/_{\tau}\subset{b}/_{\omega}$.
@@
There exist ${c},{d}\in A$, such that ${c}\in{b}/_{\omega}\setminus{b}/_{\rho}$ and ${d}\in{b}/_{\omega}\setminus{b}/_{\tau}$. We now distinguish three cases.
Case 1: $({c},{d})\in\rho$. Then by the definition of c{c}, $({b},{d})\not\in\rho$. By the definition of d{d}, $({b},{d})\not\in\tau$. Thence $({b},{d})\not\in\rho\cup\tau$.
Case 2: $({c},{d})\in\tau$. Then by the definition of d{d}, $({b},{c})\not\in\tau$. By the definition of c{c}, $({b},{c})\not\in\rho$. Consequently $({b},{c})\not\in\rho\cup\tau$.
Case 3: $({c},{d})\not\in\rho\cup\tau$. Then by the definition of c{c} and d{d}, $({c},{d})\in\omega$, and hence $({c},{d})\in\omega\setminus(\rho\cup\tau)$.
In all three cases we have ${b}/_{\rho}\cup{b}/_{\tau}\subset{b}/_{\omega}$. ∎
@@
Let $A$ be a set, ${b}\in A$, and let $\rho,\tau,\omega$ be equivalence relations on $A$ such that ${b}/_{\rho}\cup{b}/_{\tau}={b}/_{\omega}$. Then ${b}/_{\rho}={b}/_{\omega}$ or ${b}/_{\tau}={b}/_{\omega}$.
@@
By contradiction, assume that ${b}/_{\rho}\neq{b}/_{\omega}$ and ${b}/_{\tau}\neq{b}/_{\omega}$. Then ${b}/_{\rho}\subset{b}/_{\omega}$ and ${b}/_{\tau}\subset{b}/_{\omega}$. Consequently, by Proposition 9.89, ${b}/_{\rho}\cup{b}/_{\tau}\subset{b}/_{\omega}$. ∎
@@
Let $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, and ${b}_{1},\ldots,{b}_{m}\in C\langle E\cup F,\emptyset\rangle$. Then there exist $t_{1},\ldots,t_{m}\in W$ such that for every $i\in\{\,1,\ldots,m\,\}$, $t_{i}/_{\Theta\langle E\cup F,\emptyset\rangle}={b}_{i}$. As GTRS $R\langle E,F\rangle$ is total, there exists a rewrite rule $\sigma(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})\dot{\rightarrow}t/_{\Theta\langle E,F\rangle}$ in $R\langle E,F\rangle$. Accordingly, by the definition of $R\langle E,F\rangle$, $\sigma(t_{1},\ldots,t_{m})\in W$ and $\sigma(t_{1},\ldots,t_{m})\mathop{\leftrightarrow}^{*}_{E}t$. Then $\sigma(t_{1},\ldots,t_{m})\mathop{\leftrightarrow}^{*}_{E\cup F}t$. For every $i\in\{\,1,\ldots,m\,\}$, $t_{i}/_{\Theta\langle E\cup F,\emptyset\rangle}\in C\langle E\cup F,\emptyset\rangle$. Thus, by the definition of $R\langle E\cup F,\emptyset\rangle$, the rewrite rule $\sigma(t_{1}/_{\Theta\langle E\cup F,\emptyset\rangle},\ldots,t_{m}/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}t/_{\Theta\langle E\cup F,\emptyset\rangle}$ is in $R\langle E\cup F,\emptyset\rangle$. ∎
@@
$(3\Rightarrow 4)$ $\mathop{\leftrightarrow}^{*}_{H}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$, hence $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$.
@@
where for every $i\in\{\,0,\ldots,n\,\}$, $q_{i}\in T_{\Sigma}$ and for every $i\in\{\,1,\ldots,n\,\}$, $l_{i}\doteq r_{i}\in E\cup F$. We now show that
@@
If $n=0$, then $s=t$ and hence $s\mathop{\leftrightarrow}^{*}_{E}t$. Let us assume that $n\geq 1$. Let $\alpha_{i}$ be a shortest word in the set $\{\,\alpha_{1},\alpha_{2},\ldots,\alpha_{n}\,\}$, where $i\in\{\,1,\ldots,n\,\}$. Then there exist $\delta\in CON_{\Sigma,1}$, $v,z\in T_{\Sigma}$ such that
  • 1.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=\mathrm{size}(E)+\mathrm{size}(F)$. Let $W=ST\langle E\cup F\rangle$. We can decide in $O(n^{3})$ time whether for each $p\in W$,
@@
By Propositon 4.35, $s\mathop{\leftrightarrow}^{*}_{E}t$. Then
@@
We now introduce the notion of GTRS $R\langle E,F\rangle$ simulating the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant ${b}\in C\langle E\cup F,\emptyset\rangle$. We study the case when GTRS $R\langle E,F\rangle$ is total. Furthermore, we introduce the concept of GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Then we describe the connection between the latter concept and the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and let ${b}\in C\langle E\cup F,\emptyset\rangle$. We say that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant ${b}\in C\langle E\cup F,\emptyset\rangle$ if Conditions 1 and 2 hold:
@@
    For every ${c}\in C\langle E\cup F,\emptyset\rangle$, if $R\langle{E\cup F},\emptyset\rangle$ reaches b{b} from a proper extension of the constant c{c}, then
    for every rewrite rule of the form $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}\in R\langle{E\cup F},\emptyset\rangle$ with $m\in{\mathbb{N}}$,
    for all ${d}_{1},\ldots,{d}_{m}\in C\langle E,F\rangle$ with ${d}_{i}\subseteq{c}_{i}$ for all $i\in\{\,1,\ldots,m\,\}$,
    there exists ${d}\in C\langle E,F\rangle$ such that the rewrite rule $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}$ is in $R\langle E,F\rangle$.
Statement 10.99.
${d}\subseteq{c}$ for d,c{d},{c} appearing in Condition 2 in Definition 10.98.
Intuitively, in Condition 2 in Definition 10.98, the rewrite rule $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}$ in $R\langle E,F\rangle$ matches the rewrite rule $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}\in R\langle{E\cup F},\emptyset\rangle$.
By direct inspection of Definition 4.30 and Definition 10.98, we get the following statement.
Statement 10.100.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ is total, and let ${b}\in C\langle E\cup F,\emptyset\rangle$. GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant ${b}\in C\langle E\cup F,\emptyset\rangle$ if and only if ${b}\in C\langle E,F\rangle$.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. We say that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if for each constant ${b}\in C\langle E\cup F,\emptyset\rangle$, GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on b{b}.
We have the following statement by direct inspection of Definition 10.101 and Statement 10.100.
Statement 10.102.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ is total. GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant ${b}\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
1. ${b}\in C\langle E,F\rangle$.
2. ${b}\in C\langle F,E\rangle$ and GTRS $R\langle F,E\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on b{b}.
By Definitions 10.98 and 10.101, and Lemma 6.71, and Definition 6.62, we have the following lemma.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant ${b}\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
1. ${b}\in C\langle E,F\rangle$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing c{c}.
2. ${b}\in C\langle F,E\rangle$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing c{c}.
By Statement 10.102, Definition 10.98, Lemma 6.71, and Definition 6.62 we have the following lemma.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that $R\langle E,F\rangle$ is total. GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant ${b}\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
1. ${b}\in C\langle E,F\rangle$.
2. ${b}\in C\langle F,E\rangle$ and for each constant ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing c{c}.
@@
The length of the first path is $1$, and it starts from the vertex $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to $_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of $_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
$f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$   and   $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$.
The length of the second path is $1$, and it starts from the vertex $f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
$f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$   and   $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
@@
We now continue Example 3.19, Example 5.52, and Example 7.73. There does not exists a path of positive length in $AUX[E;F]$. There exists one cycle in $AUX[E;F]$. It is of length $1$, and it starts from the vertex $_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to $_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of $_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
$_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and
$_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{false}$, $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
Observe that neither Condition 1 in nor Condition 2 in Lemma 10.103 holds for the vertex $_{\Theta\langle E\cup F,\emptyset\rangle}$. Thus by Lemma 10.103, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
Intuitively, assume that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a vertex ${d}\in C\langle E\cup F,\emptyset\rangle$ of $AUX[E;F]$. Consider a term $\delta[{b}_{1},\ldots,{b}_{k}]$ for some $\delta\in CON_{\Sigma,k}$, ${b}_{1},\ldots,{b}_{k}\in C\langle{E\cup F},\emptyset\rangle$ and a sequence of reductions
(i)   $\delta[{b}_{1},\ldots,{b}_{k}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}{d}$.
Furthermore, consider any constants ${c}_{1},\ldots,{c}_{k}\in C\langle E,F\rangle$ such that ${c}_{j}\subseteq{b}_{j}$ for every $j\in\{\,1,\ldots,k\,\}$. We substitute ${c}_{j}$ for ${b}_{j}$ in $\delta[{b}_{1},\ldots,{b}_{k}]$, in this way, we get the term $\delta[{c}_{1},\ldots,{c}_{k}]$. Since GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$, we construct a fitting sequence of reductions
(ii)   $\delta[{c}_{1},\ldots,{c}_{k}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}{e}$
for some ${e}\in C\langle E,F\rangle$. We obtain (ii) following step by step the reductions in (i). For each reduction of (i), we apply a fitting rule at the same position in (ii). Then for each reduction step, for the two terms $s$ and $t$ obtained in (i) and (ii), respectively, each constant ${c}\in C\langle E,F\rangle$ appearing in $t$ corresponds to the constant ${b}\in C\langle E\cup F,\emptyset\rangle$ appearing at the same position in $s$. Moreover c{c}, as congruence class, is subset of the corresponding constant b{b}. Finally, constant e{e} corresponds to constant d{d}, and ${e}\subseteq{d}$. Formally, we state and prove the following result.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant ${d}\in C\langle E\cup F,\emptyset\rangle$. Let Conditions (a) and (b) hold.
@@
    $\delta_{n}[{b}_{n1},\ldots,{b}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[{b}_{(n-1)1},\ldots,{b}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-2}[{b}_{(n-2)1},\ldots,{b}_{(n-2)k_{n-2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[{b}_{11},\ldots,{b}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{d}$,
    where $n\geq 1$, for every $i\in\{\,1,\ldots,n\,\}$, $\delta_{i}\in CON_{\Sigma,{k_{i}}}$, $k_{i}\in{\mathbb{N}}$, and ${b}_{i1},\ldots,{b}_{i{k_{i}}}\in C\langle{E\cup F},\emptyset\rangle$.
@@
    For every $j\in\{\,1,\ldots,k_{n}\,\}$, ${c}_{nj}\in C\langle E,F\rangle$ and ${c}_{nj}\subseteq{b}_{nj}$.
Then for any $i\in\{\,1,\ldots,n-1\,\}$, $j\in\{\,1,\ldots,k_{i}\,\}$, there exist ${c}_{ij}\in C\langle E,F\rangle$ with ${c}_{ij}\subseteq{b}_{ij}$, and there exists ${e}\in C\langle E,F\rangle$ with ${e}\subseteq{d}$ such that the following sequence of reductions holds:
$\delta_{n}[{c}_{n_{1}},\ldots,{c}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[{c}_{(n-1)1},\ldots,{c}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-2}[{c}_{(n-2)1},\ldots,{c}_{(n-2)k_{n-2}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\newline \delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}{e}$.